OpenStudy (anonymous):

how to prove sin(2A) = 2sinA

OpenStudy (dumbcow):

sin(2a) = 2sinacosa

OpenStudy (dumbcow):

uses angle sum identity \[\sin(a+b) = \sin a \cos b + \sin b \cos a\] in case of "2A" a=b

OpenStudy (masumanwar):

sin(a+b) formula write for b=a